OpenStudy (anonymous):

what I like to do is make fractions out of our numbers: 1 44 ,1 10 ,1 31 and add them up to equal some arbitrary number. 1 44 +1 10 +1 31 =N then I start clearing fractions and reducing as I go

OpenStudy (amistre64):

right, does that make sense?

OpenStudy (anonymous):

yes keep going

OpenStudy (amistre64):

how do we clear a fraction then? we multiply the top by the bottom so that it becomes a rather useless form of "1"

OpenStudy (amistre64):

1* anything doesnt change the value does it

OpenStudy (anonymous):

dont you cross multiply

OpenStudy (amistre64):

so lets start with multiplying it all by 44 \[44\left(\frac{1}{44}+\frac{1}{10}+\frac{1}{31}=N\right)\] \[\frac{1(44)}{44}+\frac{1(44)}{10}+\frac{1(44)}{31}=(44)N\] and reduce \[\frac{1}{1}+\frac{22}{5}+\frac{44}{31}=44N\] next we would clear off that "5" by doing the exact same process

OpenStudy (amistre64):

\[5\left(\frac{1}{1}+\frac{22}{5}+\frac{44}{31}=44N\right)\] \[\frac{1(5)}{1}+\frac{22(5)}{5}+\frac{44(5)}{31}=44(5)N\] reduce \[\frac{5}{1}+\frac{22}{1}+\frac{44(5)}{31}=220N\] the left side is only there to help the process, so that 44(5) doesnt reduce so we can leave it there ... next is to clear that 31 off in the same process

OpenStudy (amistre64):

\[31\left(\frac{5}{1}+\frac{22}{1}+\frac{44(5)}{31}=220N\right)\] \[\frac{5(31)}{1}+\frac{22(31)}{1}+\frac{44(5)(31)}{31}=220(31)N\] and reduce \[\frac{5(31)}{1}+\frac{22(31)}{1}+\frac{44(5)}{1}=6820N\] andwith all the fractions cleared we are left with our LCM; its the number stuck to our N 6820

OpenStudy (anonymous):

so i would go 1(31)/31 + 1(31)/10 + 1(31)/31

OpenStudy (amistre64):

no, you would go as I did above. I have no idea where you got that from

OpenStudy (amistre64):

its similar yes, so i think your on the right track at least :)

OpenStudy (anonymous):

44(1/31 +1/10 + 1/44)

OpenStudy (amistre64):

You have to include an arbitrary number to collect it all on the left with

OpenStudy (amistre64):

otherwise, your just throwing out the baby with the bathwater

OpenStudy (anonymous):

ok show me what im leaving out please

OpenStudy (amistre64):

44(1/31 +1/10 + 1/44 = N)

OpenStudy (amistre64):

its that =N part; it subtle, but important. You dont have to name it N of course. name it LCM if you want, but stick it in there

OpenStudy (anonymous):

44(1/31 +1/10 + 1/44)=N

OpenStudy (amistre64):

almost, we need to include it in the process; so put it inside of the whole multiplication paranthesis

OpenStudy (amistre64):

44(1/31 +1/10 + 1/44 = N )

OpenStudy (anonymous):

OK i see i was trying to leave it out by accednt

OpenStudy (amistre64):

now distribute the 44 thru all the terms inside the paranthesis and then simplify your results in order to go at it again

OpenStudy (anonymous):

1 (44)/10 +1(44)/10 +1(44)/31

OpenStudy (anonymous):

sorry 1(44)/44

OpenStudy (amistre64):

soo close, but lets use what we created instead of making up something else. 44(1/31 +1/10 + 1/44 = N ) 44/31 + 44/10 + 44/44 = 44 N 44/31 + 22/5 + 1/1 = 44 N

OpenStudy (anonymous):

so i did it wrong

OpenStudy (amistre64):

you did it almost right :)

OpenStudy (amistre64):

you catched your 44/44, but you left off that = N part.

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

now its important to know that this is just a method that I use. If there is an easier method for you, by all means use it

OpenStudy (anonymous):

im going to tell you im having to reteach myself it has been that long

OpenStudy (amistre64):

i started college 20 years after high school :)

OpenStudy (anonymous):

now that leads us with what

OpenStudy (anonymous):

44 as the LCM

OpenStudy (amistre64):

no, 44 is just the first round of 3 rounds to finish 44(1/31 +1/10 + 1/44 = N ) 44/31 + 44/10 + 44/44 = 44 N 44/31 + 22/5 + 1/1 = 44 N we have 2 more fractions to clear out

OpenStudy (amistre64):

since 5 and 31 have nothing at all in common; we could shorten this to: 44(5)(31) and be done with it

OpenStudy (anonymous):

mulitply it it together

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

6820

OpenStudy (amistre64):

thats it :)

OpenStudy (anonymous):

now the last part GCF

OpenStudy (amistre64):

gonna have to post that one up again on the left side; i have to run

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

OK